OpenStudy (albert0898):

WILL GIVE MEDAL! Please explain! 52. Given that (x+2) and (x-1) are factors of the quadratic expression below, what are the values of a and b? x^2 + (a+2)x + a + b a b F. -4 5 G. -3 1 H. -3 5 J. -1 3 K. -1 -1

OpenStudy (albert0898):

\[x^2 + (a+2)x + a + b\]

OpenStudy (whpalmer4):

What do you get if you multiply \((x+2)(x-1)\)

OpenStudy (albert0898):

\[x^2+x-2\]

OpenStudy (whpalmer4):

Okay, so let's mix and match. \[x^2 + (a+2)x + a + b\]\[x^2 + x -2\] What do \(a,b\) have to be so that \((a+2)x = x\) and \(a+b = -2\)?

OpenStudy (albert0898):

I haven't a clue...

OpenStudy (whpalmer4):

Oh, come on. \[(a+2) x = x\]Solve that for \(a\)

OpenStudy (albert0898):

ax + 2x = x ax = x - 2x a = 1 - 2

OpenStudy (whpalmer4):

so \(a =\)

OpenStudy (albert0898):

-1

OpenStudy (whpalmer4):

right. now how about \(a+b =-2\) given that \(a = -1\)?

OpenStudy (albert0898):

b = 1

OpenStudy (albert0898):

OH NOW I GET IT!!!!

OpenStudy (albert0898):

Because x^2 - sum x + product = 0

OpenStudy (whpalmer4):

that's also a good way of looking at it, yes

OpenStudy (albert0898):

One of those problems that you just have to write out and not do in your head... duh oh Thank you very much!

OpenStudy (whpalmer4):

when you have factor a hairy polynomial, you can use that insight that the constant term is always a product of the constant terms of the factors to make educated guesses as to what might be a factor

OpenStudy (whpalmer4):

Yeah, there are a lot of problems like this where you just write down one thing and write down another thing like it and say "what do I have to do to make these match?"

OpenStudy (albert0898):

Understood, thanks again!

OpenStudy (whpalmer4):

you bet! I always like watching the "oh, I see!" moments :-)